OpenStudy (anonymous):

the length of a rectangle is fixed at 26 cm. what widths will make the perimeter greater than 98cm

OpenStudy (anonymous):

2x+52=98, 2x=46, x=23 or greater

OpenStudy (anonymous):

since the length is fixed at 26 cm and perimeter = 2l + 2w = 98 we can fill in 26 for l so we have 52 + 2w = 98 and from there we can subtract 52 so we get 2w = 46, divide by two and you get 23. or greater

OpenStudy (anonymous):

thanks. that helps

OpenStudy (anonymous):

that helps too :P _____________________________________________________________^